# DirectMultiStep: Direct Route Generation for Multistep Retrosynthesis

### v2

@@ Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a series of transformer-based models, utilizing mixture of experts approach, that directly generate multistep synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones. Our models can accommodate specific conditions such as the desired number of steps and starting materials, with the top-performing DMS-Flex (Duo) surpassing state-of-the-art methods on the PaRoutes dataset with a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. While the current suboptimal diversity of the training set may impact performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
@@ 2.2 Route Representation
[…]
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (blue), (optional) starting material (red), and the number of steps are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multistep synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (blue), (optional) starting material (red), and the predicted synthesis tree with structures of all molecules.
2.3 Evaluation Metrics and Post-Processing
Evaluation is performed on n1 and n5 (10 000 routes each) evaluation sets from PaRoutes [44], which have been created by selecting the most diverse routes from the USPTO. Fig. 2 shows the distribution of route lengths in training set, n1, and n5. The distribution in both n1 and n5 is shifted towards longer routes; for example, the fraction of routes with 4 steps or more is 15% for training set, 27% for n1 and 51% for n5.
@@ 2.3 Evaluation Metrics and Post-Processing
The structure of target compound, starting material (optionally, the deepest leaf of the correct route is chosen during evaluation), and number of steps is passed to the transformer encoder. A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder. Within beam search, during the prediction of each token the cumulative log likelihoods are normalized by the sequence length at the time the new token is sampled. This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability. After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set (following procedure of PaRoutes taken to be the set of all leaves in n1 or n5, and absence of repetitions (which can be present as permutations of other routes). Remaining routes are compared against the true (experimental) route. The reproduction of experimental routes is presented as Top-$K$ accuracy, i.e. presence of correct route ranked $K$ or lower. Importantly, the Top-$K$ accuracy is evaluated against all 10 000 targets, not just the ones for which routes are found.
2.4 Models Presented
As a baseline, we train a conventional transformer-based sequence-to-sequence model with 6 encoder and 6 decoder layers (DMS-Flash) with structures of starting materials provided during training. To achieve comparable performance without starting material as a hint, a significantly larger model with 12 encoder and 36 decoder layers is needed (DMS-Deep). We also present the first, to our knowledge, application of mixture-of-experts (MoE [46, 47, 48]) architecture to a chemical problem. To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts). A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them. Output of each active expert is added with weights determined by the router. Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal. We present two MoE models, the DMS-Flex (trained with SM) and DMS-Wide (trained without SM). The relevant architectural details of models are given in Table 1.
DMS Model Encoder Decoder Hidden FF Total Active Experts
Flash 6 6 256 3 1 1 1
Explorer 6 6 256 3 3 2 2
Explorer XL 6 24 256 3 3 2 2
Flex (Mono) 6 6 256 3 3 2 1
Flex (Duo) 6 6 256 3 3 2 2
Deep 12 36 256 3 1 1 1
Wide 12 12 256 3 3 2 2
@@ 3 Results and Discussion
3.1 Top-K Accuracy on Evaluation Sets
3.1.1 Vanilla Transformers
Top-$K$ accuracy is shown in Table 2 for set-n1 and Table 3 for set-n5. The smallest DMS-Flash model (which receives structure of SM during generation) achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.4x on n1 and 3.7x on n5. To achieve comparable performance without providing the structure of the SM, a significantly larger model DMS-Deep is needed. This establishes a trade-off between size of the model (and thus generation speed) and the amount of information it receives. The limitations of the necessity to provide the structure of the SM are discussed in Section 3.1.3. Quite remarkably, a Flash model trained with structures of SM may generate routes even when SMs are not provided (Tables 2-3). While that results in a 20% drop in Top-1 accuracy, it is significantly better than a model with the same architecture that is trained without SM (Tables S6-S7). In other words, providing structure of a correct starting material helps the model learn better and improves performance even if the starting material is not used during evaluation.
@@ 3.1.1 Vanilla Transformers
Remarkably, the Top-2 accuracy of DMS-Flash and the Top-4 accuracy of DMS-Deep exceed the Top-10 accuracy of all search methods that utilize iterative application of single-step models, highlighting the advantages of the multistep-first approach. To understand this performance difference, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node, as all reactants leading to a compound are generated at the same time. Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig. 1 is predicted before the immediate right child. This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
A few words on the relevance of Top-$K$ accuracy as the primary evaluation metric are in order. One might question whether the number of targets for which valid (i.e., having valid SMILES) routes are found (or solvability rate if all SMs are in a given stock set) and the average length of routes predicted should be considered. It is essential to realize that even if you ensure that every single-step transformation corresponds to a known reaction by template matching, the route may still be experimentally not feasible simply because atoms that are unchanged (i.e., the ones ignored by the template) contribute to the energies of the reacting orbitals through inductive and resonance effects, thereby changing the energetic favorability of the reaction. As a result, short of performing highly-accurate ab initio calculations of energies (after corresponding geometry optimization) with all relevant solvation effects and considering thermal corrections, there is no computational metric that would assess experimental feasibility of the predicted route. In addition, one could argue that the ability to discover an experimentally verified route is a prerequisite for finding shorter, but still feasible, routes. Therefore, Top-$K$ accuracy is the closest proxy to route quality. Additionally, the ability to predict correct routes with better ranks is a significant improvement because it reduces the number of candidates to be passed through the aforementioned highly expensive energy calculations. Finally, one could argue that single-step methods do not provide inherent ranking of candidates, so the performance of search methods may depend on the choice of ranking metric. However, that is yet another benefit of our approach: there is no need in devising another algorithm or model for route ranking, the beam probabilities are sufficiently accurate.
@@
The superior performance of DMS models does not incur additional computational cost. A search with a beam size of 50 takes 3-16 seconds (depending on the number of steps) with a DMS-Flash or a DMS-Flex model on a single GPU (NVIDIA M4000), which is comparable to the time required to find the first successful route using previous methods (7-50 seconds [44]).
@@
3.1.2 MoE Transformers
The DMS-Flex (Flexible), a mixture-of-experts model trained with two out of three experts active, provides the largest performance improvement when evaluated with two experts active (Duo mode): 2.5x increase in Top-1 on n1 and 3.9x on n5. Evaluating DMS-Flex with one expert active (Mono) mode worsens the performance by 30%; given that the number of valid routes (Tables S1-S2) does not change significantly, one could consider the Mono mode as a source of diversity in predictions. Notably, even though the DMS-Flex is trained with SM, it can be evaluated without providing SM (empty padding tokens are fed into the encoder instead) which still results in 34% Top-1 accuracy and 47+Top-10. Evaluating DMS-Flex without providing the number of steps as an input drops performance significantly, but a different model with the same architecture trained without number of steps as an input, DMS-Explorer (Fig. S1 shows the distribution of route lengths predicted by Explorer relative to the length of an experimental route), results in only a 8% drop in accuracy.
Using a larger DMS-Wide model trained without SM provides comparable performance to that of DMS-Deep. An important advantage of DMS-Wide is that tokens passed through different active experts can be processed on a GPU in parallel, which enables faster predictions. In addition, DMS-Wide has fewer layers, so fewer consecutive calculations have to be performed; in practice, doubling the number of encoder layers results in doubling of generation time. One might also conclude that similar performance of DMS-Deep and DMS-Wide, which have comparable parameter counts, implies that the expressive power mostly depends on the number of parameters, and not on the extra non-linearities provided by larger number of layers.
Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S7), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 85% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect the accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig. 3). Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S8-S11.
@@ 3.1.3 Use Cases for Different DMS Model Variants
The requirement to provide the starting material structure as input to the model might, in theory, seem as a limitation, limiting the relevance of DMS-Flash/Flex models. However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time. Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 7.
In addition, one can use more computationally expensive DMS-Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS-Flash/Flex to generate suggestions for each starting material of interest.
To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer-XL once. Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input. We believe better performance could be attainable in the future once bigger datasets become available. Finding ways to improve performance even with current dataset constitutes an important area of further research.
3.2 Separation of Long Routes into Shorter Routes

### v3

@@ Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a series of transformer-based models, that leverage a mixture of experts approach to directly generate multistep synthetic routes as a single string, conditionally predicting each transformation based on all preceding ones. Our DMS Explorer XL model, which requires only target compounds as input, outperforms state-of-the-art methods on the PaRoutes dataset with 1.9x and 3.1x improvements in Top-1 accuracy on the n1 and n5 test sets, respectively. Providing additional information, such as the desired number of steps and starting materials, enables both a reduction in model size and an increase in accuracy, highlighting the benefits of incorporating more constraints into the prediction process. The top-performing DMS-Flex (Duo) model scores 25-50% higher on Top-1 and Top-10 accuracies for both n1 and n5 sets. Additionally, our models successfully predict routes for FDA-approved drugs not included in the training data, demonstrating strong generalization capabilities. While the limited diversity of the training set may affect performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
@@ 2.1 Definitions
[…]
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (blue), starting material (red, optional), and the number of steps (optionally) are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multistep synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (blue), starting material (red, optional), and the predicted synthesis tree with structures of all molecules.
@@ 2.2 Route Representation
2.3 Route Generation and Post-Processing
The structure of target compound, starting material (optionally), and number of steps (optionally) is passed to the transformer encoder. A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder. Within beam search, during the prediction of each token the cumulative log likelihoods are normalized by the sequence length at the time the new token is sampled. This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability. After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set, and absence of repetitions (which can be present as permutations of other routes). Remaining routes are compared against the true (experimental) route. The reproduction of experimental routes is presented as Top-$K$ accuracy, i.e. presence of correct route ranked $K$ or lower.
@@ 3 Results and Discussion
3.1 Overview
In this work, we present DMS Explorer XL, a mixture-of-experts based model (50M parameters), that predicts full multistep retrosynthetic route given the structure of the target compound. Our model shows comparable search performance on ChemBL-5000[46] and significantly improves Top-1 accuracy on n1 and n5 evaluation datasets from PaRoutes[44]. We find that if one is willing to provide extra information to the encoder, such as desired route length and (or) starting material, the model size can be reduced and simultaneously both Top-1 and Top-10 accuracy can be increased significantly. We present a family of such models (model cards are in Sec. 5.2) and discuss their use cases.
3.2 Search Performance
First, we report the solved rate on ChemBL-5000 [46] (Table 1). Our baseline DMS Explorer XL (50M) shows comparable performance to that of a tree search on predictions from a single-step model (AiZynthFinder [46]). A slightly smaller DMS Wide (40M) model that was trained with desired route length as an extra parameter to the encoder shows significantly better performance. Despite 7x more calls to the model (with different route length parameters), the average run time per target is only 3x higher. Notably, a 4x smaller DMS Flash (10M) model trained with both route length and structure of the starting material, can be evaluated on new targets even without the structure of the starting material with a minor (3-5%) decrease in the solved rate, but at the same cost as the DMS Explorer XL model.
Table 1: Search performance for DirectMultiStep models on ChemBL-5000 [46]
Methoda Stock Setb Solved Rate Run Time (s)c
DMS Explorer XL Buyables 68.66% 26.9
AiZynthFinder [46] In-House + Enamine 75.24% NA
DMS Explorer XL eMolecules Screening 75.58% 26.9
DMS Flashd,e Buyables 86.66% 23.1
DMS Flashd,e eMolecules Screening 88.54% 23.1
DMS Wided Buyables 91.20% 75.4
DMS Wided eMolecules Screening 91.94% 75.4
a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16).
b eMolecules Screening is from Chen et al. [39] and contains 23.1M screening compounds from eMolecules as of 2019 (2019-11-01). Buyables is from Roh et al. [47] and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork. In-House + Enamine is from Genheden et al. [46] and is not publicly available.
c Averages over all targets. NA indicates that the run time is not available in the original publication.
d Uses step counts from 2 to 8 (total of 7 DMS model runs).
e SMs not provided for DMS-Flash.
Next, we report the solved rate on n1 and n5 evaluation sets (each contains 10 000 routes) in Table 2. Following the procedure of PaRoutes [44] the starting material stock set is taken to be the set of all leaves in n1 or n5. DMS Explorer XL performs noticeably worse than MCTS or Retro*, but scores better than DFPN on n5. The number and the distribution of solved targets before enforcing the stock set is shown in Tables S1-S2 and Figures S1-S2. Other DMS models that accept desired route length (Wide) and also the structure of the starting material (Flash) solve more targets, but still underperform Retro*.
Unfortunately, the training set of PaRoutes contains routes that include 47 of the targets from the commonly used USPTO-190 evaluation subset, which makes direct comparison with existing models unfair. For completeness, we report solved rate on all 190 and 143 (excluding the 47) targets in Table S3.
Table 2: Search performance for DirectMultiStep models on n1 and n5 evaluation sets
Methoda Evaluation Set Solved Targets First Solutionb Search Time (s)b
DMS-Explorer-XL n1 8008 NA 14.7
DMS-Wide n1 8089 NA 6.0
DFPN [44] n1 8475 43.0 347.3
DMS-Flash n1 8814 NA 2.0
MCTS [44] n1 9714 8.6 303.3
Retro* [44] n1 9726 7.0 300.7
DFPN [44] n5 7382 53.2 297.9
DMS-Explorer-XL n5 7904 NA 16.3
DMS-Wide n5 7950 NA 7.4
DMS-Flash n5 8646 NA 2.5
MCTS [44] n5 9676 11.7 365.7
Retro* [44] n5 9703 10.5 349.2
a All DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16).
b Averages over all targets.
3.3 Top-K Accuracy (Route Quality)
An attentive reader might notice the absence of a measure of a feasibility of proposed transformations during the discussion of solved rate above. One might argue that the number of solved targets only matters if the proposed transformations to the set of commercially available materials can be performed experimentally. In the absence of a cheap ab initio tool to predict experimental feasibility of the proposed transformation, following Genheden et al. [44] and Liu et al. [43], we utilize Top-K accuracy on reproducibility of experimentally verified routes as the proxy for route quality.
@@ 3.3 Top-K Accuracy (Route Quality)
We report Top-$K$ accuracy on n1 and n5 sets, each containing 10 000 routes. Notably, both sets were constructed to maximize internal diversity (see the PaRoutes paper [44] for the procedure), and as seen in Fig. 2, differ significantly from the training set in route length distribution. The DMS Explorer XL model (only target required) shows significant improvement on Top-1 accuracy for both n1 (1.8x, Table 3) and n5 (3.1x, Table 4) sets, while matching (n1) or improving (n5) performance on Top-10 accuracy.
@@
We attribute this improvement to the multistep-first nature of our approach. To understand the distinction from existing approaches, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node: all reactants leading to a compound are generated at the same time. Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig. 1 is predicted before the immediate right child. This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
@@
The superior performance of DMS models does not incur additional computational cost. A search with a beam size of 50 takes 1-40 seconds (depending on the complexity of the target and model size) on a single NVIDIA A100 GPU, which is comparable to the time required to find the first successful route using previous methods (7-50 seconds) [44].
3.3.1 Input information and model size trade-off
A unifying pattern of results in Tables 3-4 is that if one is willing to provide more information to the model, model size can be reduced and performance can be improved at the same time. For example, providing the desired route length reduces model size by 20% and improves accuracy by 2-4% (DMS Deep). The distribution of generated route lengths compared to the length of the experimental route is shown in Fig. S3-S4. Alternatively, providing the structure of the starting material (SM) allows to reduce the parameter count in half, and the resulting DMS Explorer (19M) model scores 4-7% better on all Top-$K$ accuracies. Providing desired route length can further half parameter count (DMS Flash is only 10M) and add another 4-5% in Top-1 and Top-10 accuracy. Notably, models that take extra information during training might still be used without that extra information, and would perform better than models of the same size trained without extra information. For example, a DMS Flash is trained with SM but might be evaluated without the structure of the SM, and the resulting performance (Tables 3-4) is 4-7% better than that of equivalent 6-layer transformer trained without SM (entry 6x3-6x3 in Tables S4-S5). These results suggest a new training strategy if one is constrained by computational resources: extra input, derived directly from training data (DMS Flash is trained with the structure of the deepest node provided as SM) might serve as a learning aid, even if not used in route generation in production.
3.3.2 Mixture of experts architecture reduces cost of generation
Route generation time depends not only on the total number of parameters, but also on how those parameters are distributed. DMS Deep (38 M) utilizes a vanilla 36-layer decoder, meaning that every token prediction has to sequentially pass through 36 attention and multi-layer perceptron blocks. We find that redistributing the model parameters from the number of layers into the width of each layer by employing mixture-of-experts [48, 49, 50] based architecture (see Sec. 5.2 for details) allows a threefold reduction in the number of layers (and thus a threefold speedup): DMS Wide (38M) needs only 12-layer decoder and performs 1-2% better than DMS Deep. Another important advantage of MoE-based models is that tokens passed through different active experts can be processed in parallel, which speeds up route generation even further.
Preliminary experiments showed that decoders with fewer than 6 layers perform significantly worse, making it impossible to train a size-comparable MoE alternative to DMS Flash. However, a 6-layer MoE trained with structures of SM (DMS Flex, 19M) similarly outperforms DMS Flash by 1-2% at no increase in average runtime given that extra parameters are processed in parallel expert blocks. Just like DMS Flash, DMS Flex can be evaluated without SM and still have better Top-1 accuracy on both n1/n5 and better Top-10 on n5 than existing methods. Interestingly, MoE-based models provide another dimension for flexibility: routes can be generated with a number of active experts different than that used during training. While DMS Flex (trained with 2 active experts) evaluated with 2 active experts (Duo mode) shows higher Top-K accuracy than if evaluated with 1 active expert (Mono mode), both result in comparable number of routes at different stages of post-processing (Tables S1-S2). As a result, one could consider predictions from Mono mode as a source of diversity.
Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S8), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
@@ 3.3.2 Mixture of experts architecture reduces cost of generation
3.3.3 Route quality as a function of route length
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 85% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect the accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S9-S12.
@@ 3.3.3 Route quality as a function of route length
3.3.4 Separation of Long Routes into Shorter Routes
@@ 3.3.5 Use Cases for Different DMS Model Variants
The requirement to provide the starting material structure as an input to the model might, in theory, seem as a limitation, limiting the relevance of DMS Flash/Flex models. However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time. Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 7. Alternatively, one can use more computationally expensive DMS Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS Flash/Flex to generate suggestions for each starting material of interest.
To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer XL once. Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input. We believe better performance could be attainable in the future once bigger datasets become available. Finding ways to improve performance even with current dataset constitutes an important area of further research.